OpenStudy (lena772):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold.

OpenStudy (lena772):

Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences.

OpenStudy (lena772):

they have same slope, what else?

OpenStudy (anonymous):

hmm......

OpenStudy (anonymous):

are you sure they have same slopes?

OpenStudy (lena772):

I'm pretty sure, I might be wrong though

OpenStudy (anonymous):

okay the slopes are equal ...

OpenStudy (anonymous):

the lines are parallel

OpenStudy (lena772):

2x + 3y = 1,500 3y=-2x+1500 y=-2/3x+500

OpenStudy (anonymous):

both the lines are parallel

OpenStudy (lena772):

That's the y value for the second month right?

OpenStudy (anonymous):

wait what is you objective?

OpenStudy (lena772):

Similarities and differences of the function for month 1 and month 2

OpenStudy (lena772):

the similarities are the same slope, and you said this also means they are parellel

OpenStudy (lena772):

the graph of the function, not the function.

OpenStudy (anonymous):

similarities : same slopes ... differences : y intercepts and x intercepts are different ...

OpenStudy (anonymous):

hmm..... what else?

OpenStudy (anonymous):

similar distances between any point b/w them

OpenStudy (lena772):

How do you calculate the x-intercept?

OpenStudy (anonymous):

take value of y=0

OpenStudy (lena772):

ok... :/

OpenStudy (lena772):

and what do you mean similar distances

OpenStudy (anonymous):

okay i think i go the similarities part 1)same slope 2)both are not parabolas ir straight lines 3)both are parallel 4)both are separated by a constant distance i dunno .. i guess this might me correct

OpenStudy (lena772):

Why are they not straight lines?

OpenStudy (anonymous):

differences: 1)one graph is higher than the other 2)x intercept is different 3)y intercept is differnt they are straight lines...they are not parabolas

OpenStudy (anonymous):

they are straight lines ... just replace that ir by ie

OpenStudy (lena772):

oh ok :)

OpenStudy (lena772):

Thanks!